OpenStudy (anonymous):

What is the percent composition by mass of hydrogen in NH4HCO3 (gram-formula= 79 grams/mole

OpenStudy (anonymous):

hi welcome to openstudy @bigbills! :) Follow these steps to solve your problem 1. You need to gather the information you need using the periodic table. You need to get the molar mass of hydrogen. 2. In the compound, count how many hydrogen atoms are present. 3. Next, you have to know the molar mass of the whole compound, which I think is already given in your question. 4. Then apply this formula using the values you gathered from steps 1-3: \( \sf Percent\ Mass\ composition= \Large{\frac{no.\ of\ H\ atoms \times molar\ mass\ of\ hydrogen}{molar\ mass\ of\ the\ compound}} \times 100\)